OpenStudy (anonymous):

John has 15 more music CD's than Michael. Together they have a total of 55 CD's. How many CD's does Michael have?

Parth (parthkohli):

Let's say that John has x cd's, then Michael has x - 15. \(\Large \color{MidnightBlue}{\Rightarrow x + (x - 15) = 55 }\) Tell me that you know how to solve for equations.

OpenStudy (anonymous):

Oh so Michael has 55 CD's

Parth (parthkohli):

nopes

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2x - 15 = 55 }\) \(\Large \color{MidnightBlue}{\Rightarrow 2x = 70 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 35 }\) Mike has x - 15 ----> 35 - 15 = 20

Parth (parthkohli):

Michael has 20 CDsss

OpenStudy (anonymous):

Ohh so Michael has 20 CD's?

Parth (parthkohli):

Yep

OpenStudy (anonymous):

Thanks(:

Parth (parthkohli):

:p